OpenStudy (joftheworld):

Quick Computing Company produces calculators. They have found that the cost, c(x), of making x calculators is a quadratic function in terms of x. The company also discovered that it costs $61 to produce 3 calculators, $165 to produce 7 calculators, and $385 to produce 12 calculators. Find the total cost of producing 6 calculators.

OpenStudy (saifoo.khan):

You are basically given three sets of values. Like an ordered pair (x, y). Here, (3, 61), (7, 165) and (12, 385) are given.

OpenStudy (joftheworld):

i know but am i supposed to plug them into three equations ?

OpenStudy (saifoo.khan):

General form is: y = ax^2 +bx + c We have to find a, b and c. Suppose y is the cost and x are the number of calculators. 61 = a(3)^2 + b(3) + c 165 = a(7)^2 + b(7) + c 185 = a(12)^2 + b(12) + c

OpenStudy (joftheworld):

so how do i solve that ? do i have to plug in more numbers?